# Jaime Fernández Fisac and Chen Lv

### v2

@@ Abstract
Causal confusion presents a considerable obstacle when applying imitation learning-based models to closed-loop simulations. Consistent with prior studies (De Haan et al., 2019; Cheng et al., 2024; Bansal et al., 2018), our findings suggest that discarding historical information from agent representations and focusing solely on their current states substantially alleviates this problem. Specifically, we employ an aggressive dropout rate (50%) for the past agent trajectories at training time and only supply the current joint state to the model during inference. As shown in Table 5, experimental results reveal that models trained and tested with history dropout surpass those using full historical states encoding across all metrics, underscoring its effectiveness in mitigating causal confusion. Specifically, we observe the baseline with full past trajectories overly accelerates or decelerates in closed-loop simulation and leads to significant high displacement errors, highlighting its copycatting behavior through kinematic history. By adopting history dropout, we achieve enhanced performance in closed-loop simulations, highlighting the significance of suitable state representation in such contexts.
@@
Table 5: Impact of History Encoding and Integrating Behavior Prior Prediction in VBD
@@
Table 6: Closed-loop Simulation Results with DDIM Sampling
@@
Collision [%] $\downarrow$ Off-road [%] $\downarrow$ ADE [$m$] $\downarrow$ Runtime [s] $\downarrow$ Collision [%] $\downarrow$ Off-road [%$\downarrow$ ADE [$m$] $\downarrow$ Runtime [s] $\downarrow$
@@
DDPM 2.50 1.96 1.42 $1.48\pm 0.005$ 1.11 1.58 1.68 $6.33\pm 0.130$
8.4 8.4 Integration of Behavior Prediction
Another key design choice in VBD is the integration of a multi-modal behavior predictor within the model architecture as a co-training task. The results shown in Table 5 indicate the inclusion of the behavior predictor significantly enhances performance in closed-loop simulation tests compared to a baseline approach lacking prediction capabilities. The integration of the predictor contributes to the refinement of the diffusion policy by facilitating more effective learning within the model’s encoder. This interaction strengthens the encoder’s capacity to extract meaningful features, which are critical for downstream tasks. Moreover, the predictor provides behavior priors for individual agents, thereby enabling flexible behavior editing and enhancing the controllability of the diffusion policy.
8.5 8.5 Accelerating Inference
@@
Employing Denoising Diffusion Implicit Models (DDIM) (Song et al., 2020a) can significantly accelerate the inference process of the VBD model while maintaining performance. To evaluate the impact of diffusion steps on simulation and runtime performance, we conduct experiments with varying numbers of diffusion steps in DDIM, and the detailed results are summarized in Table 6. The runtime of each inference step is measured on a system equipped with an AMD 7950X CPU and an NVIDIA RTX 4090 GPU. The results reveal that the runtime grows linearly with the number of diffusion steps. Notably, employing 5 or 10 diffusion steps achieves a favorable balance between generation quality and real-time performance (with a feasible replan frequency of 1 Hz). Especially when combined with collision cost guidance, we can significantly reduce collision rate and enhance the quality of generated behaviors. Specifically, using 5 diffusion steps in DDIM guided sampling yields a notable reduction in collision rates and improves scenario realism without compromising computational efficiency. Moreover, increasing the number of diffusion steps results in diminishing returns for collision avoidance and a negative impact on the log divergence (ADE) metric. This suggests the performance and efficiency of VBD can be further enhanced with DDIM sampling to balance computational efficiency and performance and retain a certain level of controllability.
9 9 Conclusions
9.1 9.1 Summary
In this work, we present Versatile Behavior Diffusion (VBD), a novel traffic scenario generation framework that integrates a diffusion generative model to produce realistic and controllable multi-agent interactions in closed-loop settings. Our approach has demonstrated state-of-the-art performance on the Waymo Sim Agents Benchmark, showcasing its capability to generate diverse, coherent, and interactive traffic scenarios that accommodate complex agent interactions and varied environmental conditions. The flexibility of VBD is highlighted through its support for controllable inference-time scenario editing and behavior guidance, allowing users to generate scenarios according to specific requirements. Our extensive evaluation and analysis reveal the effectiveness of VBD in handling a wide range of traffic simulation applications, including those involving safety-critical scenarios. Moreover, we provide valuable insights into successful training and inference strategies for diffusion-based traffic scenario generation models, shedding light on best practices and common pitfalls.
9.2 9.2 Future Work
While VBD has achieved impressive results, there are several avenues for future research and development. One important direction involves optimizing the model’s runtime efficiency to facilitate seamless integration with autonomous vehicle planning and training pipelines. Another exciting area of exploration is the automation of scenario generation using large language models (LLMs) that can learn to create context-dependent guidance objectives and prior distributions. Furthermore, expanding the scope of VBD to incorporate diverse traffic datasets, such as those from highway environments, will be crucial for improving generation quality and scalability. By pursuing these research directions, we envision VBD becoming an important tool for advancing traffic simulation and autonomous driving system development.
Appendix A Appendix
We provide an extended analysis to show that one interpretation of score-based or diffusion-based generative model under imitation learning setting is learning the gradient descent step of a particular optimal control solver based on previous results from Du and Mordatch (2019); Liu et al. (2022); Chi et al. (2023).
Consider a dataset $\mathbb{D}$ with scenario triplets $\mathcal{S}:=({\mathbf{x}},\mathbf{u},\mathbf{c})$ sampled independently from an unknown distribution $p$, whose probability density function can be factorized the as $p(\mathcal{S})=p(\mathbf{u}|\mathbf{c})p(\mathbf{c})$. In particular, we are interested in modeling the conditional probability $p(\mathbf{u}|\mathbf{c})$ of the joint control sequence $\mathbf{u}$ w.r.tthe scene context $\mathbf{c}$ by $p_{\theta}(\mathbf{u}|\mathbf{c})$, from which we can sample (high probability) $\mathbf{u}$ and implement realistic trajectories ${\mathbf{x}}$ with known dynamics ${f}$.
Under Maximum Entropy IRL (Ziebart et al., 2008) formulation, $p_{\theta}(\mathbf{u}|\mathbf{c})$ can be structured as the Boltzmann distribution of an optimization objective:
where $Z_{\theta}$ is the partition function (normalizing factor). Equation 26 resembles the Energy-Based Models (EBM), which can be trained through the estimation of maximum likelihood (LeCun et al., 2006; Song and Kingma, 2021).
Corollary 1.
The maximum likelihood estimation of $\theta$ can be obtained by maximizing conditional log-likelihood of the dataset $\mathbb{D}$:
$\theta=\operatorname*{arg\,max}_{\hat{\theta}}{\mathbb{E}}_{\mathcal{S}\sim\mathbb{D}}[\log p_{\hat{\theta}}(\mathbf{u}|\mathbf{c})].$ (27)
Proof.
Assuming we want to model the unknown data distribution $p$ by $p_{\theta,\phi}(\mathcal{S})=p_{\theta}({\mathbf{x}},\mathbf{u}|\mathbf{c})p_{\phi}(\mathbf{c})$. To estimate both $\theta$ and $\phi$, we can maximize the log-likelihood as:
$\max_{\hat{\theta},\hat{\phi}}{\mathbb{E}}_{\mathcal{S}\sim\mathbb{D}}[\log p_{\hat{\theta},\hat{\phi}}]=\max_{\hat{\theta},\hat{\phi}}{\mathbb{E}}_{\mathcal{S}\sim\mathbb{D}}[\log p_{\hat{\theta}}({\mathbf{x}},\mathbf{u}|\mathbf{c})+\log p_{\hat{\phi}}(\mathbf{c})]$
$=\max_{\hat{\theta},\hat{\phi}}{\mathbb{E}}_{\mathcal{S}\sim\mathbb{D}}[\log p_{\hat{\theta}}({\mathbf{x}},\mathbf{u}|\mathbf{c})]+\max_{\hat{\theta},\hat{\phi}}{\mathbb{E}}_{\mathcal{S}\sim\mathbb{D}}[\log p_{\hat{\phi}}(\mathbf{c})]$
$=\max_{\hat{\theta}}{\mathbb{E}}_{\mathcal{S}\sim\mathbb{D}}[\log p_{\hat{\theta}}({\mathbf{x}},\mathbf{u}|\mathbf{c})]+\max_{\hat{\phi}}{\mathbb{E}}_{\mathcal{S}\sim\mathbb{D}}[\log p_{\hat{\phi}}(\mathbf{c})]$
Therefore, we can separate the problem and learn the parameter $\theta$ of conditional distribution via maximizing log-likelihood independent of learning $p_{\phi}(\mathbf{c})$. ∎∎
Ideally, we can train a score function $\mathbf{s}_{\theta}(\mathbf{u}|\mathbf{c})$ by score-matching (Hyvärinen and Dayan, 2005; Vincent, 2011; Song and Ermon, 2019; Song et al., 2020b) to approximate the gradient of $\log p(\mathbf{u}|\mathbf{c})$ w.r.t the control sequence (our random variable of interest):
[…]
With the score function, $\nabla_{\mathbf{u}}\mathcal{J}_{\theta}$ is naturally obtained and can be directly used for gradient descent. However, since the dataset contains mostly near-optimal scenarios, the gradient estimation in suboptimal regions of the action space (away from demonstration data) may be inaccurate. Given a context $\mathbf{c}$, if we initialize an arbitrary $\mathbf{u}$ far away from the minimizer of $\mathcal{J}$, errors in the score function may lead to suboptimal generation. To overcome this issue, we follow Song et al. (2020c) to diffuse the control $\mathbf{u}$ through a forward stochastic differential equation (SDE):
$d\tilde{\mathbf{u}}=h(\tilde{\mathbf{u}},k)dk+g(k)d\mathbf{w},$ (29)
where $\mathbf{w}$ is the wiener process and $k\in[0,K]$ is the continuous diffusion step. This process allows $p$ to gradually transform into noised distributions $p_{k}$ until it becomes a known distribution $p_{K}=\pi$. While the probability density function of $p_{k}$ is unknown, we can sample a noised control $\tilde{\mathbf{u}}\sim p_{k}$ by first sampling $\mathbf{u}\sim p$ and perturbing with $p_{k}(\tilde{\mathbf{u}}|\mathbf{u})$ through SDE. This allows us to train a step-conditioned score function $\mathbf{s}_{\theta}(\tilde{\mathbf{u}}|\mathbf{c},k)$ to approximate $\nabla_{\tilde{\mathbf{u}}}\log p_{k}(\tilde{\mathbf{u}})$ by:
$\begin{split}\theta=\operatorname*{arg\,max}_{\hat{\theta}}\ {\mathbb{E}}_{\mathcal{S}\sim p,k\sim{\mathcal{U}}(0,K)}{\mathbb{E}}_{\tilde{\mathbf{u}}\sim p_{k}(\cdot|\mathbf{u})}\\ \left[\lambda(k)\lVert\nabla_{\tilde{\mathbf{u}}}\log p_{k}(\tilde{\mathbf{u}}|\mathbf{u})-\mathbf{s}_{\hat{\theta}}(\tilde{\mathbf{u}}|\mathbf{c},k)\rVert\right],\end{split}$ (30)
where $\lambda(k)$ is a positive weighting function. At inference time, we can generate scenarios by first sampling $\tilde{\mathbf{u}}$ from the known distribution $\pi$ and iterating through the reverse SDE (Anderson, 1982):
$d\tilde{\mathbf{u}}=\left[h(\tilde{\mathbf{u}},k)-g^{2}(k)\mathbf{s}_{\theta}(\tilde{\mathbf{u}}|\mathbf{c},k)\right]dk+g(k)d\mathbf{\bar{w}}.$ (31)
To connect the score-based model with IRL and EBM, we can view the forward diffusion as uplifting original data distribution into a higher-dimensional space. By injecting noise, we achieve good coverage over the entire action space in the final step $K$ so that $\mathbf{s}_{\theta}(\tilde{\mathbf{u}}|\mathbf{c},K)$ are well defined for random $\tilde{\mathbf{u}}$. Sampling through reverse SDE can be interpreted as stochastic gradient descent towards high-probability regions with a fixed descent direction along the diffusion step, analogous to the direct shooting method in optimal control. We note that at low noise level $k$, as $p_{k}$ is close to the original data distribution $p$, $\mathbf{s}_{\theta}(\mathbf{u}|\mathbf{c},k)\approx-\nabla_{\mathbf{u}}\mathcal{J}_{\theta}({\mathbf{x}},\mathbf{u};\mathbf{c})$, which recovers our learning objective. Therefore, the generative modeling of scenarios can be viewed as an implicit solution of IRL by learning the gradient steps of trajectory optimization and solving the optimal control problem through sampling.

### v3

@@ IX-B Influence of History Encoding
Causal confusion presents a considerable obstacle when applying imitation learning-based models to closed-loop simulations. Consistent with prior studies [10, 8, 1], our findings suggest that discarding historical information from agent representations and focusing solely on their current states substantially alleviates this problem. Specifically, we employ an aggressive dropout rate (50%) for the past agent trajectories at training time and only supply the current joint state to the model during inference. As shown in Table IV, experimental results reveal that models trained and tested with history dropout surpass those using full historical states encoding across all metrics, underscoring its effectiveness in mitigating causal confusion. Specifically, we observe that the baseline with full past trajectories overly accelerates or decelerates in closed-loop simulation and leads to significantly high displacement errors, highlighting its copying behavior through kinematic history. By adopting history dropout, we achieve enhanced performance in closed-loop simulations, highlighting the significance of suitable state representation.
@@
TABLE IV: Impact of History Encoding and Integrating Behavior Prior Prediction in VBD
@@
IX-C Integration of Behavior Prediction
Another key design choice in VBD is the integration of a multi-modal behavior predictor within the model architecture as a co-training task. The results shown in Table IV indicate that the inclusion of the behavior predictor significantly enhances performance in closed-loop simulation tests compared to a baseline approach lacking prediction capabilities. The integration of the predictor contributes to the refinement of the diffusion policy by facilitating more effective learning within the model’s encoder. This interaction strengthens the encoder’s capacity to extract meaningful features, which are critical for downstream tasks. Moreover, the predictor provides behavior priors for individual agents, thereby enabling behavior editing and enhancing the controllability of the diffusion policy.
TABLE V: Closed-loop Simulation Results with DDIM Sampling
@@ IX-C Integration of Behavior Prediction
Collision [%] $\downarrow$ Off-road [%] $\downarrow$ ADE [$m$] $\downarrow$ Runtime [s] $\downarrow$ Collision [%] $\downarrow$ Off-road $\downarrow$ ADE [$m$] $\downarrow$ Runtime [s] $\downarrow$
@@
DDPM 2.50 1.96 1.42 $48\pm 0.005 1.1$1 1.58 1.68 $6.33\pm 0.130$
IX-D Noise Scheduling Design
Traditional cosine scheduling tends to favor low noise variance regions during the initial diffusion steps, potentially allowing the model to bypass proper scene representation learning by relying on shortcuts through less noisy inputs. Consequently, when encountering out-of-distribution actions during testing, the model struggles to adapt and instead produces degraded output by directly passing the actions through the final denoising steps. This insight motivated us to develop a logarithmic noise schedule, presented in Eq. 14, which maintains a suitable signal-to-noise ratio (SNR) throughout the diffusion process, ensuring effective action denoising that is dependent on scene context. Additionally, experimental results in Fig. 11 validate the efficacy of our proposed approach, showcasing improved trajectory diffusion performance under the new noise schedule. By adopting the log scheduling strategy, VBD successfully generates diverse scenarios without compromising generation quality.
The result in Fig. 11 reveals that the generation quality is satisfactory after just the first denoising step, especially regarding collision rate, off-road rate, ADE, and FDE. Further denoising steps seem to degrade these metrics. One possible explanation is that traffic behavior diffusion is heavily influenced by scene context; hence, the initial denoising step can yield plausible outcomes that adhere to both agent interactions and map constraints. Nevertheless, subsequent denoising steps play a vital role in striking a balance between fidelity and diversity. With increasing numbers of steps, we notice a decline in the minADE and minFDE metrics, suggesting that one of the generated samples draws closer to the ground truth. Average ADE and FDE metrics have increased, highlighting greater sample diversity. This outcome demonstrates that additional denoising steps enable the model to explore a wider range of feasible outcomes while preserving high-quality trajectories. Furthermore, these extra steps are crucial for facilitating control and guidance over agent behaviors, particularly in applications such as traffic scenario generation.
IX-E Accelerating Inference
@@
Employing Denoising Diffusion Implicit Models (DDIM) [48] can significantly accelerate the inference process of the VBD model while maintaining performance. To evaluate the impact of diffusion steps on simulation and runtime performance, we conduct experiments with varying numbers of diffusion steps in DDIM, and the detailed results are summarized in Table V. The runtime of each inference step is measured on a system equipped with an AMD 7950X CPU and an NVIDIA RTX 4090 GPU. The results reveal that the runtime grows linearly with the number of diffusion steps. Notably, employing 5 or 10 diffusion steps achieves a favorable balance between generation quality and real-time performance (with a feasible replan frequency of 1 Hz). Especially when combined with collision cost guidance, we can significantly reduce the collision rate and enhance the quality of the generated behaviors. Specifically, using 5 diffusion steps in DDIM guided sampling yields a notable reduction in collision rates and improves scenario realism without compromising computational efficiency. Moreover, increasing the number of diffusion steps results in diminishing returns for collision avoidance and a negative impact on the log divergence (ADE) metric. This suggests that the performance and efficiency of VBD can be further enhanced with DDIM sampling to balance computational efficiency and performance while retaining a certain level of controllability.
X Conclusions
We present Versatile Behavior Diffusion (VBD), a traffic scenario generation framework that leverages diffusion models to produce realistic, diverse, and controllable multi-agent interactions in closed-loop simulation. Our method demonstrates strong performance in multi-agent behavior simulation and effectively captures complex interaction patterns. A key strength of VBD lies in its flexible inference-time control, which enables user-guided scenario editing and behavior conditioning to meet specific requirements. Through extensive experiments, we show that VBD is effective across a broad range of traffic simulation tasks, including safety-critical scenarios. Several directions remain for future work. A key next step is integrating VBD into AV planning and testing pipelines to enable systematic evaluation and improvement of planning algorithms under generated scenarios, including rare and adversarial cases. Enhancing runtime efficiency will also be critical for real-time deployment in AV simulation and training systems. Additionally, incorporating large language models (LLMs) for context-aware scenario generation and extending the framework to diverse traffic domains, such as highways or urban mixed-autonomy settings, are promising avenues for further development.